### v5

# Deep Compression Autoencoder for Efficient High-Resolution Diffusion Models

Abstract
We present Deep Compression Autoencoder (DC-AE), a new family of autoencoders for accelerating high-resolution diffusion models. Existing autoencoders have demonstrated impressive results at a moderate spatial compression ratio (e.g., 8), but fail to maintain satisfactory reconstruction accuracy for high spatial compression ratios (e.g., 64). We address this challenge by introducing two key techniques: (1) Residual Autoencoding, where we design our models to learn residuals based on the space-to-channel transformed features to alleviate the optimization difficulty of high spatial-compression autoencoders; (2) Decoupled High-Resolution Adaptation, an efficient decoupled three-phase training strategy for mitigating the generalization penalty of high spatial-compression autoencoders. With these designs, we improve the autoencoder’s spatial compression ratio up to 128 while maintaining the reconstruction quality. Applying our DC-AE to latent diffusion models, we achieve significant speedup without accuracy drop. For example, on ImageNet , our DC-AE provides 19.1 inference speedup and 17.9 training speedup on H100 GPU for UViT-H while achieving a better FID, compared with the widely used SD-VAE-f8 autoencoder.
1 Introduction
Latent diffusion models (Rombach et al., 2022) have emerged as a leading framework and demonstrated great success in image synthesis (Labs, 2024; Esser et al., 2024). They employ an autoencoder to project the images to the latent space to reduce the cost of diffusion models. For example, the predominantly adopted solution in current latent diffusion models (Rombach et al., 2022; Labs, 2024; Esser et al., 2024; Chen et al., 2024b; a) is to use an autoencoder with a spatial compression ratio of 8 (denoted as f8), which converts images of spatial size to latent features of spatial size . This spatial compression ratio is satisfactory for low-resolution image synthesis (e.g., ). However, for high-resolution image synthesis (e.g., ), further increasing the spatial compression ratio is critical, especially for diffusion transformer models (Peebles & Xie, 2023; Bao et al., 2023) that have quadratic computational complexity to the number of tokens.
The current common practice for further reducing the spatial size is downsampling on the diffusion model side. For example, in diffusion transformer models (Peebles & Xie, 2023; Bao et al., 2023), this is achieved by using a patch embedding layer with patch size that compresses the latent features to tokens. In contrast, little effort has been made on the autoencoder side. The main bottleneck hindering the employment of high spatial-compression autoencoders is the reconstruction accuracy drop. For example, Figure 2 (a) shows the reconstruction results of SD-VAE (Rombach et al., 2022) on ImageNet with different spatial compression ratios. We can see that the rFID (reconstruction FID) degrades from 0.90 to 28.3 if switching from f8 to f64.
This work presents Deep Compression Autoencoder (DC-AE), a new family of high spatial-compression autoencoders for efficient high-resolution image synthesis. By analyzing the underlying source of the accuracy degradation between high spatial-compression and low spatial-compression autoencoders, we find high spatial-compression autoencoders are more difficult to optimize (Section 3.1) and suffer from the generalization penalty across resolutions (Figure 3 b). To this end, we introduce two key techniques to address these two challenges. First, we propose Residual Autoencoding (Figure 4) to alleviate the optimization difficulty of high spatial-compression autoencoders. It introduces extra non-parametric shortcuts to the autoencoder to let the neural network modules learn residuals based on the space-to-channel operation. Second, we propose Decoupled High-Resolution Adaptation (Figure 6) to tackle the other challenge. It introduces a high-resolution latent adaptation phase and a low-resolution local refinement phase to avoid the generalization penalty while maintaining a low training cost.
With these techniques, we increase the spatial compression ratio of autoencoders to 32, 64, and 128 while maintaining good reconstruction accuracy (Table 2). The diffusion models can fully focus on the denoising task with our DC-AE taking over the whole token compression task, which delivers better image generation results than prior approaches (Table 3). For example, replacing SD-VAE-f8 with our DC-AE-f64, we achieve 17.9 higher H100 training throughput and 19.1 higher H100 inference throughput on UViT-H (Bao et al., 2023) while improving the ImageNet FID from 3.55 to 3.01. We summarize our contributions as follows:
We analyze the challenges of increasing the spatial compression ratio of autoencoders and provide insights into how to address these challenges.
We propose Residual Autoencoding and Decoupled High-Resolution Adaptation that effectively improve the reconstruction accuracy of high spatial-compression autoencoders, making their reconstruction accuracy feasible for use in latent diffusion models.
We build DC-AE, a new family of autoencoders based on our techniques. It delivers significant training and inference speedup for diffusion models compared with prior autoencoders.
2 Related Work
Autoencoder for Diffusion Models.
Training and evaluating diffusion models directly in high-resolution pixel space results in prohibitive computational costs. To address this issue, Rombach et al. (2022) proposes latent diffusion models that operate in a compressed latent space produced by pretrained autoencoders. The proposed autoencoder with spatial compression ratio and latent channels has been widely adopted in subsequent works (Peebles & Xie, 2023; Bao et al., 2023). Since then, follow-up works mainly focus on enhancing the reconstruction accuracy of the f8 autoencoder by increasing the number of latent channels (Esser et al., 2024; Dai et al., 2023; Labs, 2024). Additionally, to improve the reconstruction quality, Zhu et al. (2023) leverages a heavier decoder and incorporates task-specific priors. In contrast to prior works, our work focuses on an orthogonal direction, increasing the spatial compression ratio of the autoencoders (e.g., f64). To the best of our knowledge, our work is the first study in this critical but underexplored direction.
Diffusion Model Acceleration.
Diffusion models have been widely used for image generation and showed impressive results (Labs, 2024; Esser et al., 2024). However, diffusion models are computationally intensive, motivating many works to accelerate diffusion models. One representative strategy is reducing the number of inference sampling steps by training-free few-step samplers (Song et al., 2021; Lu et al., 2022a; b; Zheng et al., 2023; Zhang & Chen, 2023; Zhang et al., 2023; Zhao et al., 2024b; Shih et al., 2024; Tang et al., 2024) or distilling-based methods (Meng et al., 2023; Salimans & Ho, 2022; Yin et al., 2024b; a; Song et al., 2023; Luo et al., 2023; Liu et al., 2023). Another representative strategy is model compression by leveraging sparsity (Li et al., 2022; Ma et al., 2024b) or quantization (He et al., 2024; Fang et al., 2024; Li et al., 2023; Zhao et al., 2024a). Designing efficient diffusion model architectures (Li et al., 2024d; Liu et al., 2024; Cai et al., 2024) or inference systems (Li et al., 2024b; Wang et al., 2024) is also an effective approach for boosting efficiency. In addition, improving the data quality (Chen et al., 2024b; a) can boost the training efficiency of diffusion models.
All these works focus on diffusion models while the autoencoder remains the same. Our work opens up a new direction for accelerating diffusion models, which can benefit both training and inference.
3 Method
In this section, we first analyze why existing high spatial-compression autoencoders (e.g., SD-VAE-f64) fail to match the accuracy of low spatial-compression autoencoders (e.g., SD-VAE-f8). Then we introduce our Deep Compression Autoencoder (DC-AE) with Residual Autoencoding and Decoupled High-Resolution Adaptation to close the accuracy gap. Finally, we discuss the applications of our DC-AE to latent diffusion models.
3.1 Motivation
We conduct ablation study experiments to get insights into the underlying source of the accuracy gap between high spatial-compression and low spatial-compression autoencoders. Specifically, we consider three settings with gradually increased spatial compression ratio, from f8 to f64.
Each time the spatial compression ratio increases, we stack additional encoder and decoder stages upon the current autoencoder. In this way, high spatial-compression autoencoders contain low spatial-compression autoencoders as sub-networks and thus have higher learning capacity.
Additionally, we increase the latent channel number to maintain the same total latent size across different settings. We can then convert the latent to a higher spatial compression ratio one by applying a space-to-channel operation (Shi et al., 2016): .
We summarize the results in Figure 3 (a, gray dash line). Even with the same total latent size and stronger learning capacity, we still observe degraded reconstruction accuracy when the spatial compression ratio increases. It demonstrates that the added encoder and decoder stages (consisting of multiple SD-VAE building blocks) work worse than a simple space-to-channel operation.
Based on this finding, we conjecture the accuracy gap comes from the model learning process: while we have good local optimums in the parameter space, the optimization difficulty hinders high spatial-compression autoencoders from reaching such local optimums.
3.2 Deep Compression Autoencoder
Residual Autoencoding.
Motivated by the analysis, we introduce Residual Autoencoding to address the accuracy gap. The general idea is depicted in Figure 4. The core difference from the conventional design is that we explicitly let neural network modules learn the downsample residuals based on the space-to-channel operation to alleviate the optimization difficulty. Different from ResNet (He et al., 2016), the residual here is not identity mapping, but space-to-channel mapping.
In practice, this is implemented by adding extra non-parametric shortcuts on the encoder’s downsample blocks and decoder’s upsample blocks (Figure 4 b, left). Specifically, for the downsample block, the non-parametric shortcut is a space-to-channel operation followed by a non-parametric channel averaging operation to match the channel number. For example, assuming the downsample block’s input feature map shape is and its output feature map shape is , then the added shortcut is:
Accordingly, for the upsample block, the non-parametric shortcut is a channel-to-space operation followed by a non-parametric channel duplicating operation:
In addition to the downsample and upsample blocks, we also change the middle stage design following the same principle (Figure 4 b, right).
Figure 3 (a) shows the comparison with and without our Residual Autoencoding on ImageNet . We can see that Residual Autoencoding effectively improves the reconstruction accuracy of high spatial-compression autoencoders.
Decoupled High-Resolution Adaptation.
Residual Autoencoding alone can address the accuracy gap when handling low-resolution images. However, when extending it to high-resolution images, we find it not sufficient. Due to the large cost of high-resolution training, the common practice for high-resolution diffusion models is directly using autoencoders trained on low-resolution images (e.g., ) (Chen et al., 2024b; a). This strategy works well for low spatial-compression autoencoders. However, high spatial-compression autoencoders suffer from a significant accuracy drop. For example, in Figure 3 (b), we can see that f64 autoencoder’s rFID degrades from 0.50 to 7.40 when generalizing from to . In contrast, the f8 autoencoder’s rFID improves from 0.51 to 0.19 under the same setting. Additionally, we also find this issue more severe when using a higher spatial compression ratio. In this work, we refer to this phenomenon as the generalization penalty of high spatial-compression autoencoders. A straightforward solution to address this issue is conducting training on high-resolution images. However, it suffers from a large training cost and unstable high-resolution GAN loss training.
We introduce Decoupled High-Resolution Adaptation to tackle this challenge. Figure 6 demonstrates the detailed training pipeline. Compared with the conventional single-phase training strategy (Rombach et al., 2022), our Decoupled High-Resolution Adaptation has two key differences.
First, we decouple the GAN loss training from the full model training and introduce a dedicated local refinement phase for the GAN loss training. In the local refinement phase (Figure 6, phase 3), we only tune the head layers of the decoder while freezing all the other layers. The intuition of this design is based on the finding that the reconstruction loss alone is sufficient for learning to reconstruct the content and semantics. Meanwhile, the GAN loss mainly improves local details and removes local artifacts (Figure 5). Achieving the same goal of local refinement, only tuning the decoder’s head layers has a lower training cost and delivers better accuracy than the full training.
Moreover, the decoupling prevents the GAN loss training from changing the latent space. This approach enables us to conduct the local refinement phase on low-resolution images without worrying about the generalization penalty. This further reduces the training cost of phase 3 and avoids the highly unstable high-resolution GAN loss training.
Second, we introduce an additional high-resolution latent adaptation phase (Figure 6, phase 2) that tunes the middle layers (i.e., encoder’s head layers and decoder’s input layers) to adapt the latent space for alleviating the generalization penalty. In our experiments, we find only tuning middle layers is sufficient for addressing this issue (Figure 3 b) while having a lower training cost than high-resolution full training (memory cost: 153.98 GB 67.81 GB)111Assuming the input resolution is and the batch size is 12. (Cai et al., 2020).
| ImageNet 512512 (Class-Conditional) | |||||
| Diffusion Model | Autoencoder | Patch Size | #Tokens | FID (w/o CFG) | FID (w/ CFG) |
| SD-VAE-f8 | 8 | 64 | 125.08 | 95.93 | |
| SD-VAE-f16 | 4 | 64 | 115.32 | 88.06 | |
| SD-VAE-f32 | 2 | 64 | 107.33 | 76.57 | |
| UViT-S [1] | DC-AE-f64 | 1 | 64 | 67.30 | 35.96 |
3.3 Application to Latent Diffusion Models
Applying our DC-AE to latent diffusion models is straightforward. The only hyperparameter to change is the patch size (Peebles & Xie, 2023). For diffusion transformer models (Peebles & Xie, 2023; Bao et al., 2023), increasing the patch size is the common approach for reducing the number of tokens. It is equivalent to first applying the space-to-channel operation to reduce the spatial size of the given latent by and then using the transformer model with a patch size of 1.
Since combining a low spatial-compression autoencoder (e.g., f8) with the space-to-channel operation can also achieve a high spatial compression ratio, a natural question is how it compares with directly reaching the target spatial compression ratio with DC-AE.
We conduct ablation study experiments and summarize the results in Table 1. We can see that directly reaching the target spatial compression ratio with the autoencoder gives the best results among all settings. In addition, we also find that shifting the spatial compression ratio from the diffusion model to the autoencoder consistently leads to better FID.
4 Experiments
| ImageNet 256256 | Latent Shape | Autoencoder | rFID | PSNR | SSIM | LPIPS |
| f32c32 | 8832 | SD-VAE [37] | 2.64 | 22.13 | 0.59 | 0.117 |
| DC-AE | 0.69 | 23.85 | 0.66 | 0.082 | ||
| f64c128 | 44128 | SD-VAE [37] | 26.65 | 18.07 | 0.41 | 0.283 |
| DC-AE | 0.81 | 23.60 | 0.65 | 0.087 | ||
| ImageNet 512512 | Latent Shape | Autoencoder | rFID | PSNR | SSIM | LPIPS |
| f64c128 | 88128 | SD-VAE [37] | 16.84 | 19.49 | 0.48 | 0.282 |
| DC-AE | 0.22 | 26.15 | 0.71 | 0.080 | ||
| f128c512 | 44512 | SD-VAE [37] | 100.74 | 15.90 | 0.40 | 0.531 |
| DC-AE | 0.23 | 25.73 | 0.70 | 0.084 | ||
| FFHQ 10241024 | Latent Shape | Autoencoder | rFID | PSNR | SSIM | LPIPS |
| f64c128 | 1616128 | SD-VAE [37] | 6.62 | 24.55 | 0.68 | 0.237 |
| DC-AE | 0.23 | 31.04 | 0.83 | 0.061 | ||
| f128c512 | 88512 | SD-VAE [37] | 179.71 | 18.11 | 0.63 | 0.585 |
| DC-AE | 0.41 | 31.18 | 0.83 | 0.062 | ||
| MapillaryVistas 20482048 | Latent Shape | Autoencoder | rFID | PSNR | SSIM | LPIPS |
| f64c128 | 3232128 | SD-VAE [37] | 7.55 | 22.37 | 0.68 | 0.262 |
| DC-AE | 0.36 | 29.57 | 0.84 | 0.075 | ||
| f128c512 | 1616512 | SD-VAE [37] | 152.09 | 17.82 | 0.67 | 0.594 |
| DC-AE | 0.38 | 29.70 | 0.84 | 0.074 |
4.1 Setups
Implementation Details.
We use a mixture of datasets to train autoencoders (baselines and DC-AE), containing ImageNet (Deng et al., 2009), SAM (Kirillov et al., 2023), MapillaryVistas (Neuhold et al., 2017), and FFHQ (Karras et al., 2019). For ImageNet experiments, we exclusively use the ImageNet training split to train autoencoders and diffusion models. The model architecture is similar to SD-VAE (Rombach et al., 2022) except for our new designs discussed in Section 3.2. In addition, we use the original autoencoders instead of the variational autoencoders for our models, as they perform the same in our experiments and the original autoencoders are simpler. We also replace transformer blocks with EfficientViT blocks (Cai et al., 2023) to make autoencoders more friendly for handling high-resolution images while maintaining similar accuracy.
For image generation experiments, we apply autoencoders to diffusion transformer models including DiT (Peebles & Xie, 2023) and UViT (Bao et al., 2023). We follow the same training settings as the original papers. Additionally, we build USiT by combining UViT (Bao et al., 2023) with the SiT sampler (Ma et al., 2024a). We consider three settings with different resolutions, including ImageNet (Deng et al., 2009) for generation, FFHQ (Karras et al., 2019) and MJHQ (Li et al., 2024a) for generation, and MapillaryVistas (Neuhold et al., 2017) for generation.
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | FID | |||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | w/o CFG | w/ CFG |
| Flux-VAE-f8 [18] | 2 | 54 | 416 | 31.7 | 56.3 | 27.35 | 8.72 | |
| Asym-VAE-f8 [54] | 2 | 54 | 424 | 31.7 | 56.2 | 11.55 | 2.95 | |
| SD-VAE-f8 [37] | 2 | 54 | 424 | 31.7 | 56.2 | 12.03 | 3.04 | |
| DC-AE-f32 | 1 | 241 | 2016 | 7.8 | 20.9 | 9.56 | 2.84 | |
| DiT-XL [35] | DC-AE-f32‡ | 1 | 241 | 2016 | 7.8 | 20.9 | 6.88 | 2.41 |
| Flux-VAE-f8 [18] | 2 | 55 | 349 | 30.4 | 54.2 | 30.91 | 12.63 | |
| Asym-VAE-f8 [54] | 2 | 55 | 351 | 30.3 | 54.1 | 11.36 | 3.51 | |
| SD-VAE-f8 [37] | 2 | 55 | 351 | 30.3 | 54.1 | 11.04 | 3.55 | |
| DC-AE-f32 | 1 | 247 | 1622 | 8.2 | 18.6 | 9.83 | 2.53 | |
| DC-AE-f64 | 1 | 984 | 6706 | 3.5 | 10.6 | 13.96 | 3.01 | |
| UViT-H [1] | DC-AE-f64† | 1 | 984 | 6706 | 3.5 | 10.6 | 12.26 | 2.66 |
| Asym-VAE-f8 [54] | 2 | 27 | 157 | 74.8 | OOM | 9.87 | 3.62 | |
| SD-VAE-f8 [37] | 2 | 27 | 157 | 74.8 | OOM | 9.73 | 3.57 | |
| DC-AE-f32 | 1 | 112 | 665 | 19.7 | 42.0 | 8.13 | 2.30 | |
| DC-AE-f64 | 1 | 450 | 2733 | 8.6 | 30.2 | 7.78 | 2.47 | |
| UViT-2B [1] | DC-AE-f64† | 1 | 450 | 2733 | 8.6 | 30.2 | 6.50 | 2.25 |
| MAGVIT-v2 [48] | - | - | - | - | - | - | 3.07 | 1.91 |
| EDM2-XXL [15] | SD-VAE-f8 [37] | - | - | - | - | - | 1.91 | 1.81 |
| MAR-L [22] | - | - | - | - | - | - | 2.74 | 1.73 |
| USiT-H | DC-AE-f32 | 1 | 247 | 1622 | 8.2 | 18.6 | 3.80 | 1.89 |
| USiT-2B | DC-AE-f32 | 1 | 112 | 665 | 19.7 | 42.0 | 2.90 | 1.72 |
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | MJHQ 512512 | |||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | FID | CLIP Score |
| SD-VAE-f8 [37] | 2 | 43 | 312 | 37.1 | 60.45 | 6.3 | 26.36 | |
| PIXART- [6] | DC-AE-f32 | 1 | 173 | 1251 | 10.4 | 23.77 | 6.1 | 26.41 |
Efficiency Profiling.
We profile the training and inference throughput on the H100 GPU with PyTorch and TensorRT respectively. The latency is measured on the 3090 GPU with batch size 2. The training memory is profiled using PyTorch, assuming a batch size of 256. We use fp16 for all cases. For simplicity, we assume the number of sampling steps is 1.
4.2 Image Compression and Reconstruction
Table 2 summarizes the results of DC-AE and SD-VAE (Rombach et al., 2022) under various settings (f represents the spatial compression ratio and c denotes the number of latent channels). DC-AE provides significant reconstruction accuracy improvements than SD-VAE for all cases. For example, on ImageNet , DC-AE improves the rFID from 16.84 to 0.22 for the f64c128 autoencoder and 100.74 to 0.23 for the f128c512 autoencoder.
In addition to the quantitative results, Figure 7 shows image reconstruction samples produced by SD-VAE and DC-AE. Reconstructed images by DC-AE demonstrate a better visual quality than SD-VAE’s reconstructed images. In particular, for the f64 and f128 autoencoders, DC-AE still maintains a good visual quality for small text and the human face.
4.3 Latent Diffusion Models
We compare DC-AE with the widely used SD-VAE-f8 autoencoder (Rombach et al., 2022) on various diffusion transformer models. For DC-AE, we always use a patch size of 1 (denoted as p1). For SD-VAE-f8, we follow the common setting and use a patch size of 2 or 4 (denoted as p2, p4). The results are summarized in Table 3, Table 4, and Figure 9.
ImageNet 512512.
As shown in Table 3, DC-AE-f32p1 consistently delivers better FID than SD-VAE-f8p2 on all diffusion transformer models. In addition, it has 4 fewer tokens than SD-VAE-f8p2, leading to 4.5 higher H100 training throughput and 4.8 higher H100 inference throughput for DiT-XL. We also observe that larger diffusion transformer models seem to benefit more from our DC-AE (Figure 9). For example, DC-AE-f64p1 has a worse FID than SD-VAE-f8p2 on UViT-S but a better FID on UViT-2B. We conjecture it is because DC-AE-f64 has a larger latent channel number than SD-VAE-f8, thus needing more model capacity (Esser et al., 2024).
Applying DC-AE to USiT models, we achieve highly competitive results compared with prior leading image generative models. For example, DC-AE-f32+USiT-2B achieves 1.72 FID on ImageNet 512512, outperforming the SOTA diffusion model EDM2-XXL and SOTA auto-regressive image generative models (MAGVIT-v2 and MAR-L).
Text-to-Image Generation.
Table 4 reports our text-to-image generation results. All models are trained for 100K iterations from scratch. Similar to prior cases, we observe DC-AE-f32p1 provides a better FID and a better CLIP Score than SD-VAE-f8p2. Figure 8 demonstrates samples generated by the diffusion models with our DC-AE, showing the capacity to synthesize high-quality images while being significantly more efficient than prior models.
5 Conclusion
We accelerate high-resolution diffusion models by designing deep compression autoencoders to reduce the number of tokens. We proposed two techniques: residual autoencoding and decoupled high-resolution adaptation to address the challenges brought by the high compression ratio. The resulting new autoencoder family DC-AE demonstrated satisfactory reconstruction accuracy with a spatial compression ratio of up to 128. DC-AE also demonstrated significant training and inference efficiency improvements when applied to latent diffusion models.
Acknowledgements
We thank NVIDIA for donating the DGX machines. We thank MIT-IBM Watson AI Lab, MIT and Amazon Science Hub, MIT AI Hardware Program, and National Science Foundation for supporting this research.
Appendix A Additional Samples
Appendix B Diffusion Model Architecture Details
In addition to existing UViT models, we scaled the model up to 1.6B parameters, with a depth of 28, a hidden dimension of 2048, and 32 heads. We denote this model as UViT-2B.
Appendix C High-Resolution Image Generation Results
| FFHQ 10241024 (Unconditional) & MJHQ 10241024 (Class-Conditional) | |||||||||
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | FFHQ FID | MJHQ FID | |||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | w/o CFG | w/o CFG | w/ CFG |
| SD3-VAE-f8 [9] | 2 | 83 | 814 | 14.2 | 41.4 | 46.28 | 109.43 | 103.02 | |
| Flux-VAE-f8 [18] | 2 | 83 | 814 | 14.2 | 41.4 | 59.15 | 143.16 | 139.06 | |
| SDXL-VAE-f8 [36] | 2 | 84 | 833 | 14.1 | 41.2 | 16.82 | 49.00 | 39.21 | |
| Asym-VAE-f8 [54] | 2 | 84 | 833 | 14.1 | 41.2 | 17.12 | 48.25 | 38.36 | |
| 2 | 84 | 833 | 14.1 | 41.2 | 16.98 | 48.05 | 38.19 | ||
| SD-VAE-f8 [37] | 4 | 470 | 5566 | 2.5 | 10.7 | 23.81 | 60.94 | 51.29 | |
| DC-AE-f32 | 1 | 475 | 5575 | 2.5 | 10.7 | 13.65 | 34.35 | 27.20 | |
| DC-AE-f32‡ | 1 | 475 | 5575 | 2.5 | 10.7 | 11.39 | 28.36 | 21.89 | |
| DiT-S [35] | DC-AE-f64 | 1 | 2085 | 25259 | 1.0 | 3.1 | 26.88 | 61.30 | 53.38 |
| MapillaryVistas 20482048 (Unconditional) | |||||||||
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | MapillaryVistas FID | ||||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | w/o CFG | ||
| SD-VAE-f8 [37] | 4 | 84 | 810 | 14.3 | 41.4 | 69.50 | |||
| DiT-S [35] | DC-AE-f64 | 1 | 459 | 5435 | 2.6 | 11.0 | 59.55 | ||
Apart from ImageNet 512512, we also test our models for higher-resolution image generation. As shown in Table 5, we have a similar finding where DC-AE-f32p1 achieves better FID than SD-VAE-f8p2 for all cases.
Appendix D Image Generation Results with Other Evaluation Metrics
| Diffusion | Patch | Inference | FID | Inception Score | Precision | Recall | CMMD | ||||||
| Model | Autoencoder | Size | Throughput | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG |
| SD3-VAE-f8 [9] | 2 | 2984 | 164.34 | 143.82 | 6.07 | 7.53 | 0.06 | 0.09 | 0.31 | 0.39 | 3.13 | 2.94 | |
| Flux-VAE-f8 [18] | 2 | 2984 | 106.07 | 84.73 | 13.39 | 17.71 | 0.28 | 0.37 | 0.39 | 0.42 | 1.90 | 1.67 | |
| SDXL-VAE-f8 [36] | 2 | 2991 | 51.03 | 26.38 | 27.58 | 56.72 | 0.57 | 0.74 | 0.58 | 0.50 | 1.35 | 1.05 | |
| Asym-VAE-f8 [54] | 2 | 2991 | 52.68 | 25.14 | 30.22 | 65.27 | 0.58 | 0.74 | 0.62 | 0.51 | 1.09 | 0.80 | |
| SD-VAE-f8 [37] | 2 | 2991 | 51.96 | 24.57 | 30.37 | 65.73 | 0.57 | 0.74 | 0.64 | 0.52 | 1.23 | 0.91 | |
| SD-VAE-f16 [37] | 2 | 12881 | 76.86 | 44.22 | 21.38 | 43.35 | 0.43 | 0.62 | 0.60 | 0.55 | 1.83 | 1.46 | |
| SD-VAE-f32 [37] | 1 | 12883 | 70.23 | 38.63 | 23.07 | 47.72 | 0.46 | 0.64 | 0.58 | 0.56 | 1.71 | 1.36 | |
| DC-AE-f32 | 1 | 12850 | 46.12 | 18.08 | 34.82 | 84.73 | 0.59 | 0.76 | 0.66 | 0.56 | 1.00 | 0.70 | |
| DC-AE-f64 | 1 | 53774 | 67.30 | 35.96 | 24.55 | 52.86 | 0.44 | 0.64 | 0.60 | 0.56 | 1.44 | 1.14 | |
| UViT-S [1] | DC-AE-f64† | 1 | 53774 | 61.84 | 30.63 | 27.28 | 61.76 | 0.47 | 0.67 | 0.63 | 0.56 | 1.35 | 1.04 |
| Flux-VAE-f8 [18] | 2 | 416 | 27.35 | 8.72 | 53.09 | 130.20 | 0.68 | 0.83 | 0.61 | 0.48 | 0.54 | 0.30 | |
| Asym-VAE-f8 [54] | 2 | 424 | 11.39 | 2.97 | 108.70 | 241.10 | 0.75 | 0.83 | 0.65 | 0.53 | 0.37 | 0.20 | |
| SD-VAE-f8 [37] | 2 | 424 | 12.03 | 3.04 | 105.25 | 240.82 | 0.75 | 0.84 | 0.64 | 0.54 | 0.43 | 0.25 | |
| DC-AE-f32 | 1 | 2016 | 9.56 | 2.84 | 117.49 | 226.98 | 0.75 | 0.82 | 0.64 | 0.55 | 0.34 | 0.22 | |
| DiT-XL [35] | DC-AE-f32‡ | 1 | 2016 | 6.88 | 2.41 | 141.07 | 263.56 | 0.76 | 0.82 | 0.63 | 0.56 | 0.29 | 0.18 |
| Flux-VAE-f8 [18] | 2 | 349 | 30.91 | 12.63 | 56.72 | 127.93 | 0.64 | 0.76 | 0.59 | 0.49 | 0.50 | 0.31 | |
| Asym-VAE-f8 [54] | 2 | 351 | 11.36 | 3.51 | 124.24 | 249.21 | 0.75 | 0.82 | 0.61 | 0.53 | 0.32 | 0.20 | |
| SD-VAE-f8 [37] | 2 | 351 | 11.04 | 3.55 | 125.08 | 250.66 | 0.75 | 0.82 | 0.61 | 0.53 | 0.39 | 0.26 | |
| DC-AE-f32 | 1 | 1622 | 9.83 | 2.53 | 121.91 | 255.07 | 0.76 | 0.83 | 0.65 | 0.54 | 0.34 | 0.20 | |
| DC-AE-f64 | 1 | 6706 | 13.96 | 3.01 | 99.20 | 229.16 | 0.73 | 0.83 | 0.64 | 0.53 | 0.50 | 0.31 | |
| UViT-H [1] | DC-AE-f64† | 1 | 6706 | 12.26 | 2.66 | 109.20 | 239.82 | 0.73 | 0.82 | 0.67 | 0.57 | 0.43 | 0.27 |
| Flux-VAE-f8 [18] | 2 | 155 | 25.03 | 10.12 | 74.04 | 161.29 | 0.67 | 0.78 | 0.58 | 0.51 | 0.38 | 0.24 | |
| Asym-VAE-f8 [54] | 2 | 157 | 9.87 | 3.62 | 131.95 | 258.63 | 0.76 | 0.83 | 0.59 | 0.52 | 0.30 | 0.19 | |
| SD-VAE-f8 [37] | 2 | 157 | 9.73 | 3.57 | 132.86 | 260.50 | 0.76 | 0.83 | 0.59 | 0.52 | 0.37 | 0.24 | |
| DC-AE-f32 | 1 | 665 | 8.13 | 2.30 | 135.44 | 272.73 | 0.76 | 0.82 | 0.66 | 0.56 | 0.30 | 0.17 | |
| DC-AE-f64 | 1 | 2733 | 7.78 | 2.47 | 138.11 | 280.49 | 0.77 | 0.84 | 0.63 | 0.54 | 0.35 | 0.22 | |
| UViT-2B [1] | DC-AE-f64† | 1 | 2733 | 6.50 | 2.25 | 152.35 | 293.45 | 0.77 | 0.83 | 0.65 | 0.56 | 0.31 | 0.19 |
| USiT-H | DC-AE-f32 | 1 | 1622 | 3.80 | 1.89 | 174.58 | 252.35 | 0.78 | 0.82 | 0.64 | 0.60 | 0.24 | 0.18 |
| USiT-2B | DC-AE-f32 | 1 | 665 | 2.90 | 1.72 | 187.68 | 248.10 | 0.79 | 0.82 | 0.63 | 0.61 | 0.21 | 0.17 |
Table 6 presents a comprehensive evaluation of different diffusion models and autoencoders on ImageNet 512512. The evaluation metrics include FID (Martin et al., 2017), inception score (IS) (Salimans et al., 2016), precision, recall (Kynkäänniemi et al., 2019), and CMMD (Jayasumana et al., 2024). Our DC-AE consistently delivers significant efficiency improvements while maintaining the generation performance under different evaluation metrics.